can you help me with this problem sirs.Thank you..
1.) For a national day parade display, 80 girls formed a large circle.The girls were equally spaced out and numbered in a clockwise direction from 1 to 80 by wearing a number tag.If Rose had tag no. 55,what was the number of the girl directly opposite to her?If Nina had tag no. 9,what was the number of the girl directly opposite to her?

Answer by Theo(13342)

a circle is 360 degrees.
there are 80 girls.
each girl is 360/80 degrees apart from the other one.
-----
this equals 4.5 degrees between each girl.
-----
the girl opposite here would be 180 degrees from her.
180 / 4.5 = 40.
if she was girl number 1, then the girl opposite her would be number 41 because 1 + 40 = 41.
-----
to see this easier, assume there were 4 girls forming the circle.
360 / 4 = 90 degrees between each of them.
the girl opposite number 1 would be 180 degrees from her. 180/90 = 2. 1+2 = 3
-----
now if she was girl number 2, then the girl opposite her would be girl number 4.
-----
if she was girl number 3, then the girl opposite her would be girl number 1 because 3 + 2 = 5 - 4 = 1
-----
you can count this one out.
girl number 3 + 1 = girl number 4 plus 1 = girl number 5 who is the same as girl number 1.
-----
the same concept applies to the circle with a larger number of girls.
-----
80 girls form a circle.
360 degrees in the circle.
4.5 degrees between each girl.
-----
the girl opposite girl number 1 would be 180/4.5 girls from her which would make her 40 girls from her. 1 + 40 = 41.
-----
the girl opposite girl number 55 would be 180/4.5 girls from her which would make her 40 girls from her. 55 + 40 = 95 - 80 = number 15.
-----
the girl opposite girl number 9 would be 180/4.5 girls from her which would make her 40 girls from her. 9 + 40 = girl number 49
-----
Answer to your question is:
the girl opposite rose who is girl number 55 is girl number 15
the girl opposite nina who is girl number 9 is girl number 49
-----
A picture of a circle with 16 girls can be found on the following website:
http://theo.x10hosting.com/
-----
click on problem number 214346. If it's not there when you look, wait 30 minutes and try again.
-----
In that picture, there are 16 girls.
circle is 360 degrees.
degrees between each girl are 360/16 = 22.5 degrees.
-----
girl opposite girl number 1 is 180 degrees from her.
180 / 22.5 = 8
girl opposite girl number 1 is 8 girls from her.
1 + 8 = 9
she is girl number 9.
the picture confirms this.
-----
girl opposite girl number 12 is 8 girls from her.
12 + 8 = 20
20 is greater than 16 which is the number of girls, so we subtract 16 from 20 to get:
20-16 = 4
girl opposite girl number 12 is girl number 4
picture confirms that.
-----